Frank had three-fourths of the problems right on a test. Lucilla had five-sixths of the
problems right, which was ten more problems right than Frank.
a) How many problems were on the test?
b) How many problems did Frank have right?
c) How many problems did Lucilla have right?

Answer by jorel1380(3719)

a)If Lucilla had 5/6ths of the answers right, and Frank had only 3/4ths, and the difference was 10, then:
5n/6 - 3n/4=10; where n is the total number of problems on the test. So:
10n/12 - 9n/12=10
n=120 questions on the test
b)3/4 x 120=90 problems right for Frank
c)5/6 x 120=100 problems right for Lucilla
☺☺☺☺